OpenStudy (anonymous):

(1, -1); parallel to y = 2/5x - 3 How do I solve this using slope-intercept?

OpenStudy (anonymous):

Slope intercept for is y=mx+b where "m" is your slope and "b" is your y intercept the problem gives you a y intercept of -1(From (1,-1)) so that is your "b" and parallel lines have the same slope so you just need to use the slope in the given line. so your "m" = 2/5 so the final equation would be y = 2/5 x - 1

OpenStudy (anonymous):

Thank you!

OpenStudy (anonymous):

No problem